OpenStudy (anonymous):

Use the five steps for problem solving. Show All Work.. The alterations store on Birch Street charges $13 to hem a pair of jeans and $14 to hem a pair of cuffed pants. The alterations store on Westland Avenue charges $11 to hem a pair of jeans and $16 to hem a pair of cuffed pants. Roxanne has 3 pair of cuffed pants that need hemmed. How many pairs of jeans will she need to get hemmed to make the alterations store on Westland Avenue a better deal than the alterations store on Birch Street? See my answer below. Need to make sure its right..

OpenStudy (anonymous):

B = total spent at Birch J = no. of Jeans W = totals spent at Westland B=(3*14)+13J W=(3*16)+11J Simultaneous equations so to work out when they would cost the same (so W=B) we add them together to get (3*14)+13J = (3*16) +11J so 42 + 13J = 48+11J so 2J = 6 therefore J = 3 So to get a better deal she would need >3 jeans, or 4 jeans or more.